OpenStudy (nicole143):

Can you show me how to write an equation of an ellipse centered at the origin, satisfying the given conditions? @mathmale

OpenStudy (nicole143):

Conditions: Vertex - \[(0, \sqrt{29} )\] Co vertex - (-5, 0)

OpenStudy (nicole143):

@mathmale @texaschic101 Help?

OpenStudy (mathmale):

Hello, Nicole! The equation of an ellipse centered at the origin is \[\frac{ x^2 }{ a^2 }+\frac{ y^2 }{ b^2 }=1\]

OpenStudy (mathmale):

If one vertex is \[(0, \sqrt{29} )\] and a co-vertex is (-5,0), then " a " represents the greater distance of these two points from the origin. That greater distance is \[ \sqrt{29}\] and " b " represents the distance of the co-vertex (horizontally) from the origin, or 5. In summary, a=Sqrt(29) and b =5. a^2 = 29 and b^2 =25 stick these 2 values into the general equation of an ellipse centered at the origin that I gave you, above. Is this ellipse horiz. or vert.?

OpenStudy (mathmale):

If vertical, then a^2 = 29 must go under y^2 in the equation; if horizontal, then a^2 = 29 must go under x^2 in the equation.

OpenStudy (nicole143):

Oh my goodness! Thank you so much! I get everything but I'm unsure on the determination of the ellipse being horiz or vert.. I forget which is which.

OpenStudy (mathmale):

Which is greater, 5 or Sqrt(29)? That will tell you whether your ellipse is vert. or horiz. :)

OpenStudy (nicole143):

Sqrt(29) is bigger. @mathmale

OpenStudy (mathmale):

Right. Is that on the y-axis or on the x-axis? 25 goes under one of the following: x^2 or y^2; Sqrt(29) goes under the other one. Which?

OpenStudy (nicole143):

25 under x and sqrt(29) under y ??

OpenStudy (mathmale):

Exactly. Nice work, Nicole!!

OpenStudy (nicole143):

Thank you again (: